The H Balmer line (rest wavelength 656.3nm) in the optical spectrum of a giant elliptical galaxy has an observed wavelength of 659.16nm. How fast is this galaxy moving away from us? Your answer should be in km/s.

Select one:

a. 1,100 km/s

b. 1,300 km/s

c. 1,200 km/s

d. 1,400 km/s

e. 1,000 km/s
Answer #1

Let , v/c = x

Where , c is speed of light.

By Doppler effect formula

(1+x)/(1-x) = (659.16/656.3)2

(1+x)/(1-x) = 1.00873

Solving we get,

x =0.004348

v/(3×105 km/s) = 0.004348

v = 1304 km/s

Option (B)